# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 7.2. Submersion metric, connection and parallel transport
For $Y\in\mathcal{M}=\mathrm{St}_{n,d}$, with $\mathfrak{q}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)=Y$, and $\xi\in T_{Y}\mathcal{M}=T_{Y}\mathrm{St}_{n,p}$, using eq. 20, the geodesics, with initial conditions $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is
(47) $\gamma(t)=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\operatorname{expm}(t\begin{bmatrix}2\alpha A&-B^{\mathsf{T}}\\ B&0\end{bmatrix})\operatorname{expm}(t\begin{bmatrix}(1-2\alpha)A&0\\ 0&0\end{bmatrix})\operatorname{I}_{n,d},$
if $\bar{\xi}=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$. Parallel transport could also be computed using eq. 39, noting in this case, $(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu^{\mathsf{T}}\xi)_{\mathfrak{a}}=\begin{bmatrix}A&0\\ 0&0\end{bmatrix}$ commutes with $\mathfrak{k}$, and the simplified condition is satisfied. The operations are $O(n^{3})$ and not efficient. The formula is dependent on $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$.
However, we can use the isometry in section 7.1. For geodesics, the isometry gives us the geodesic formula $\gamma(t)=F\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}(t)$, where $\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}$ is the geodesic on $\mathrm{St}_{d+k,d}$ starting at $\operatorname{I}_{d+k,d}$ with initial velocity $\begin{bmatrix}A\\ R\end{bmatrix}$. We lift $\gamma_{\mathfrak{F}\mathcal{M}_{\mathcal{S}}}$ to $\mathrm{SO}(d+k)$ starting at $\operatorname{I}_{d+k}$, with initial horizontal velocity $\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$. The result is the known geodesic formula in [5, 31, 19]
@@
The parallel transport equation eq. 39 could be used for tangent vectors of $\mathcal{M}_{\mathcal{S}}$. Since parallel transports are linear in initial vectors, for the complete picture, we will need a formula for tangent vectors to $\mathrm{St}_{n,d}$ normal to $T_{Y}\mathcal{M}_{\mathcal{S}}$. We have
@@
For a geodesic $\gamma(t)$ with $\gamma(0)=Y\in\mathcal{M}=\mathrm{St}_{n,p},\dot{\gamma}(0)=\xi\in T_{Y}\mathcal{M}$, and $\mathcal{S}$ is the vector subspace of $\mathbb{R}^{n\times d}$ of matrices with columns in the span of columns of $Y$ and $\xi$ as in eq. 42. If $\eta\in T_{Y}\mathrm{St}_{n,d}$ satisfies both $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}\xi=0$, or equivalently, $\eta$ is normal to $\mathcal{S}$ in the metric eq. 44, then with $A=Y^{\mathsf{T}}\xi$, the parallel transport of $\eta$ along $\gamma$ in that metric is given by
@@
First, we show the condition $\eta$ is normal to $\mathcal{S}$ is equivalent to $Y^{\mathsf{T}}\eta=0$ and $\xi^{\mathsf{T}}\eta=0$. We need to show $\langle\eta,YA_{1}+QR_{1}\rangle_{\vec{\beta}}=0$ for all anti symmetric $A_{1}$ and $R_{1}\in\mathbb{R}^{k\times d}$ means $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}Q=0$. But $\langle\eta,YA_{1}\rangle_{\vec{\beta}}=\alpha\operatorname{Tr}\eta^{\mathsf{T}}YA_{1}$ and $\langle\eta,QR_{1}\rangle_{\vec{\beta}}=\operatorname{Tr}\eta^{\mathsf{T}}QR_{1}$ and $\eta^{\mathsf{T}}Y$ is antisymmetric, hence $\eta^{\mathsf{T}}Y=0,\eta^{\mathsf{T}}Q=0$. The reverse direction is clear.
@@
and assuming eq. 49, $\dot{\Delta}=(1-\alpha)\eta e^{t(1-\alpha)A}A$. Write $\gamma=\begin{bmatrix}Y|Q\end{bmatrix}C_{1},\dot{\gamma}=\begin{bmatrix}Y|Q\end{bmatrix}C_{2}$ then $\Delta^{\mathsf{T}}\gamma$ and $\Delta^{\mathsf{T}}\dot{\gamma}$ simplifies to
$\begin{gathered}\Delta^{\mathsf{T}}\gamma=(e^{-t(1-\alpha)A})\eta^{\mathsf{T}}\begin{bmatrix}Y|Q\end{bmatrix}C_{1}=0,\\ \Delta^{\mathsf{T}}\dot{\gamma}=(e^{-t(1-\alpha)A})\eta^{\mathsf{T}}\begin{bmatrix}Y|Q\end{bmatrix}C_{2}=0\end{gathered}$
by the initial condition. Thus, $\Gamma(\dot{\gamma},\Delta)=(1-\alpha)\Delta\dot{\gamma}^{\mathsf{T}}\gamma$ and
$\begin{gathered}\dot{\Delta}+\Gamma(\dot{\gamma},\Delta)=\dot{\Delta}+(1-\alpha)\Delta\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}-A&R^{\mathsf{T}}\\ -R&0\end{bmatrix}\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}\\ =\dot{\Delta}-(1-\alpha)\Delta A=0\end{gathered}$
which is satisfied by $\Delta(t)$ in eq. 49. ∎
@@
Let $\eta\in T_{Y}\mathrm{St}_{n,d}$ be another tangent vector at $X$. Then, the parallel transport $\Delta=\Delta(t)$ of $\eta$ along the geodesic $\gamma(t)$ with $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is given by
@@
Here, $\mathrm{P}=\mathrm{P}_{AR}$ is the operator on $\mathcal{F}:=\mathrm{Skew}_{p}\times\mathbb{R}^{k\times p}\subset\mathbb{R}^{(p+k)\times p}$, mapping $w=\begin{bmatrix}w_{a}\\ w_{r}\end{bmatrix}\in\mathcal{F}$ to $\mathrm{P}(w)$ as follows
(52) $\mathrm{P}(w)=\mathrm{P}_{AR}(w):=\begin{bmatrix}((4\alpha-1)w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}}\\ \alpha(w_{r}A-Rw_{a})\end{bmatrix}.$
@@
We can verify $\eta_{\mathcal{S}}$ is in $\mathcal{S}$ and $Y^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$, $Q^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=Q^{\mathsf{T}}\eta-Q^{\mathsf{T}}\eta=0$, hence $\xi^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$. Thus, the theorem follows from eq. 49 applied to $\eta_{\mathcal{S}^{o}}$ and the isometry in section 7.1 applied to $\eta_{\mathcal{S}}$, and section 5. It remains to derive $\mathrm{P}$ from eq. 41. With $\mathtt{b}=\begin{bmatrix}\mathtt{b}_{A}&-\mathtt{b}_{R}^{\mathsf{T}}\\ \mathtt{b}_{R}&0\end{bmatrix}$, $\mathtt{a}=\begin{bmatrix}\mathtt{a}_{A}&-\mathtt{a}_{R}^{\mathsf{T}}\\ \mathtt{a}_{R}&0\end{bmatrix}$ then $[\mathtt{b},\mathtt{a}]_{\mathfrak{m}}\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{b}_{A},\mathtt{a}_{A}]-\mathtt{b}_{R}^{\mathsf{T}}\mathtt{a}_{R}+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R}\\ \mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A}\end{bmatrix}$, $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{a}_{A},\mathtt{b}_{A}]\\ -\mathtt{b}_{R}\mathtt{a}_{A}\end{bmatrix}$, and $[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{b}_{A},\mathtt{a}_{A}]\\ -\mathtt{a}_{R}\mathtt{b}_{A}\end{bmatrix}$. Thus,
@@
$\begin{gathered}\frac{1}{2}\left([\mathtt{b}_{A},\mathtt{a}_{A}]+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R}-\mathtt{b}_{R}^{\mathsf{T}}\mathtt{a}_{R}\right)-(1-2\alpha)[\mathtt{b}_{A},\mathtt{a}_{A}]=((4\alpha-1)\mathtt{b}_{A}\mathtt{a}_{A}+\mathtt{a}_{R}^{\mathsf{T}}\mathtt{b}_{R})_{\mathrm{skew}},\\ \frac{1}{2}(\mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A})+\frac{1-2\alpha}{2}(\mathtt{a}_{R}\mathtt{b}_{A}-\mathtt{b}_{R}\mathtt{a}_{A})=\alpha(\mathtt{b}_{R}\mathtt{a}_{A}-\mathtt{a}_{R}\mathtt{b}_{A}).\end{gathered}$
giving us the top and bottom rows of $\mathrm{P}_{\mathtt{a}}\mathtt{b}$, with $A=\mathtt{a}_{A},R=\mathtt{a}_{R},w_{r}=\mathtt{b}_{R},w_{a}=\mathtt{b}_{A}$. ∎
8. Flag manifolds

### v2

@@ 8.2. Submersion metric, connection and parallel transport
For $Y\in\mathcal{M}=\mathrm{St}_{n,d}$, with $\Theta(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)=Y$, and $\xi\in T_{Y}\mathcal{M}=T_{Y}\mathrm{St}_{n,p}$, using eq. 28, the geodesics, with initial conditions $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is
(57) $\gamma(t)=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\operatorname{expm}(t\begin{bmatrix}2\alpha A&-B^{\mathsf{T}}\\ B&0\end{bmatrix})\operatorname{expm}(t\begin{bmatrix}(1-2\alpha)A&0\\ 0&0\end{bmatrix})\operatorname{I}_{n,d}$
if $\bar{\xi}=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$. Parallel transport could also be computed using eq. 50 as $\mathcal{K}$ commutes with $\mathcal{A}$. The operations are $O(n^{3})$ and not efficient.
However, we can use the isometry in section 8.1. For geodesics, the isometry gives us the geodesic formula $\gamma(t)=F^{\mathsf{T}}\gamma_{\mathsf{F}\mathcal{M}_{\mathcal{S}}}(t)$, where $\gamma_{\mathsf{F}\mathcal{M}_{\mathcal{S}}}$ is the geodesic on $\mathrm{St}_{d+k,d}$ starting at $\operatorname{I}_{d+k,d}$ with the initial velocity $\begin{bmatrix}A\\ R\end{bmatrix}$. We lift $\gamma_{\mathsf{F}\mathcal{M}_{\mathcal{S}}}$ to $\gamma_{\mathsf{F}\mathcal{M}_{\mathcal{S}}}^{\mathrm{SO}(d+k)}\in\mathrm{SO}(d+k)$ starting at $\operatorname{I}_{d+k}$, with the initial horizontal velocity $\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$. The resulting expression $F^{\mathsf{T}}(\gamma^{\mathrm{SO}(d+k)}_{\mathsf{F}\mathcal{M}_{\mathcal{S}}}(t)\operatorname{I}_{d+k,d})$ is the known geodesic formula in [7, 46, 29]
@@
The parallel transport equation eq. 50 could be used for tangent vectors of $\mathcal{M}_{\mathcal{S}}$. Since parallel transports are linear in initial vectors, for the complete picture, we will need a transport formula for tangent vectors to $\mathrm{St}_{n,d}$ normal to $T_{Y}\mathcal{M}_{\mathcal{S}}$. We have
@@
For a geodesic $\gamma(t)$ with $\gamma(0)=Y\in\mathcal{M}=\mathrm{St}_{n,p},\dot{\gamma}(0)=\xi\in T_{Y}\mathcal{M}$, let $\mathcal{S}$ be the vector subspace of $\mathbb{R}^{n\times d}$ of matrices with columns in the span of columns of $Y$ and $\xi$ as in eq. 52. If $\eta\in T_{Y}\mathrm{St}_{n,d}$ satisfies both $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}\xi=0$, or equivalently, $\eta$ is normal to $\mathcal{S}$ in the metric eq. 54, then with $A=Y^{\mathsf{T}}\xi$, the parallel transport of $\eta$ along $\gamma$ in that metric is given by
@@
First, we show the condition $\eta$ is normal to $\mathcal{S}$ is equivalent to $Y^{\mathsf{T}}\eta=0$ and $\xi^{\mathsf{T}}\eta=0$. We need to show $\langle\eta,YA_{1}+QR_{1}\rangle_{\vec{\beta}}=0$ for all antisymmetric $A_{1}$ and $R_{1}\in\mathbb{R}^{k\times d}$ means $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}Q=0$. But $\langle\eta,YA_{1}\rangle_{\vec{\beta}}=\alpha\operatorname{Tr}\eta^{\mathsf{T}}YA_{1}$ and $\langle\eta,QR_{1}\rangle_{\vec{\beta}}=\operatorname{Tr}\eta^{\mathsf{T}}QR_{1}$ and $\eta^{\mathsf{T}}Y$ is antisymmetric. Since the Frobenius form is nondegenerate on the space of antisymmetric matrices and on $\mathbb{R}^{k\times d}$, we have $\eta^{\mathsf{T}}Y=0$ and $\eta^{\mathsf{T}}Q=0$. The reverse direction is clear.
@@
Since $\gamma(t)$ and $\dot{\gamma(t)}$ stay in $\mathcal{S}$, they are are normal to $\eta$. If $\Delta$ is given by eq. 59 then $\Delta^{\mathsf{T}}\gamma=\Delta^{\mathsf{T}}\dot{\gamma}=0$ since for $S\in\mathcal{S}$, $\Delta^{\mathsf{T}}S=\operatorname{expm}(-t(1-\alpha)A)\eta^{\mathsf{T}}S=0$. Substitute in eq. 56, we get $\Gamma(\dot{\gamma},\Delta)=(1-\alpha)\Delta\dot{\gamma}^{\mathsf{T}}\gamma$ and
$\begin{gathered}\dot{\Delta}+\Gamma(\dot{\gamma},\Delta)=\dot{\Delta}+(1-\alpha)\Delta\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}^{\mathsf{T}}\begin{bmatrix}-A&R^{\mathsf{T}}\\ -R&0\end{bmatrix}\begin{bmatrix}\operatorname{expm}((1-2\alpha)tA)\\ 0_{k,d}\end{bmatrix}\\ =\dot{\Delta}-(1-\alpha)\Delta e^{-(1-2\alpha)tA}Ae^{(1-2\alpha)tA}=(1-\alpha)\eta e^{t(1-\alpha)A}A-(1-\alpha)\Delta A=0\end{gathered}$
verifying the transport equation. ∎
@@
Let $\eta\in T_{Y}\mathrm{St}_{n,d}$ be another tangent vector at $Y$. Then, the parallel transport $\Delta=\Delta(t)$ of $\eta$ along the geodesic $\gamma(t)$ with $\gamma(0)=Y,\dot{\gamma}(0)=\xi$ is given by
@@
Here, $\mathrm{P}=\mathrm{P}_{AR}$ is the operator on $\mathcal{W}:=\mathrm{Skew}_{p}\times\mathbb{R}^{k\times p}\subset\mathbb{R}^{(p+k)\times p}$ that maps $\mathtt{W}=\begin{bmatrix}\mathtt{W}_{a}\\ \mathtt{W}_{r}\end{bmatrix}\in\mathcal{W}$ to $\mathrm{P}(\mathtt{W})$ as follows
(62) $\mathrm{P}(\mathtt{W})=\mathrm{P}_{AR}(\mathtt{W}):=\begin{bmatrix}((4\alpha-1)\mathtt{W}_{a}A+R^{\mathsf{T}}\mathtt{W}_{r})_{\mathrm{skew}}\\ \alpha(\mathtt{W}_{r}A-R\mathtt{W}_{a})\end{bmatrix}$.
@@
We can verify $\eta_{\mathcal{S}}$ is in $\mathcal{S}$ and $Y^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$, $Q^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=Q^{\mathsf{T}}\eta-Q^{\mathsf{T}}\eta=0$, hence $\xi^{\mathsf{T}}\eta_{\mathcal{S}^{o}}=0$. Thus, the theorem follows from lemma 5 applied to $\eta_{\mathcal{S}^{o}}$ and the isometry in section 8.1 applied to $\eta_{\mathcal{S}}$, and section 6. It remains to derive $\mathrm{P}$ from eq. 51. With $\mathtt{B}=\begin{bmatrix}\mathtt{B}_{A}&-\mathtt{B}_{R}^{\mathsf{T}}\\ \mathtt{B}_{R}&0\end{bmatrix}$, $\mathtt{V}=\begin{bmatrix}\mathtt{V}_{A}&-\mathtt{V}_{R}^{\mathsf{T}}\\ \mathtt{V}_{R}&0\end{bmatrix}$ then $\mathcal{P}_{\mathfrak{m}}[\mathtt{B},\mathtt{V}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{B}_{A},\mathtt{V}_{A}]-\mathtt{B}_{R}^{\mathsf{T}}\mathtt{V}_{R}+\mathtt{V}_{R}^{\mathsf{T}}\mathtt{B}_{R}\\ \mathtt{B}_{R}\mathtt{V}_{A}-\mathtt{V}_{R}\mathtt{B}_{A}\end{bmatrix}$, $\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{V}_{A},\mathtt{B}_{A}]\\ -\mathtt{B}_{R}\mathtt{V}_{A}\end{bmatrix}$, and $\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}]\operatorname{I}_{d+k,d}=\begin{bmatrix}[\mathtt{B}_{A},\mathtt{V}_{A}]\\ -\mathtt{V}_{R}\mathtt{B}_{A}\end{bmatrix}$. Thus,
@@
$\begin{gathered}\frac{1}{2}\left([\mathtt{B}_{A},\mathtt{V}_{A}]+\mathtt{V}_{R}^{\mathsf{T}}\mathtt{B}_{R}-\mathtt{B}_{R}^{\mathsf{T}}\mathtt{V}_{R}\right)-(1-2\alpha)[\mathtt{B}_{A},\mathtt{V}_{A}]=((4\alpha-1)\mathtt{B}_{A}\mathtt{V}_{A}+\mathtt{V}_{R}^{\mathsf{T}}\mathtt{B}_{R})_{\mathrm{skew}},\\ \frac{1}{2}(\mathtt{B}_{R}\mathtt{V}_{A}-\mathtt{V}_{R}\mathtt{B}_{A})+\frac{1-2\alpha}{2}(\mathtt{V}_{R}\mathtt{B}_{A}-\mathtt{B}_{R}\mathtt{V}_{A})=\alpha(\mathtt{B}_{R}\mathtt{V}_{A}-\mathtt{V}_{R}\mathtt{B}_{A})\end{gathered}$
giving us the top and bottom rows of $\mathrm{P}_{\mathtt{V}}\mathtt{B}$, with $A=\mathtt{V}_{A},R=\mathtt{V}_{R},\mathtt{W}_{r}=\mathtt{B}_{R},\mathtt{W}_{a}=\mathtt{B}_{A}$. ∎
Remark 2.
The transport formula can be evaluated with cost $O(C_{1}nd^{2}+C_{2}td^{3})$ with $n>d$ because $Q,A,R$ could be computed with cost $O(nd^{2})$ [7, Corollary 2.2], the matrix exponential is of cost $O(d^{3})$ (or more precisely $O(d^{3}\log_{2}t))$, the exponential action is of cost $O(td^{3})$, and the matrix multiplications are $O(nd^{2})$. Thus, the time complexity is $O(C_{1}nd^{2}+C_{2}td^{3})$. As mentioned, for small $t$, the first $O(nd^{2})$ term dominates the cost, whereas, for large $t$, the term $O(td^{3})$ will dominate. We can say the algorithm is $O(nd^{2})$ for fixed $t$. Since Stiefel manifolds are compact, there is an upper bound for distance between two points. Thus, there is an upper bound for reasonable times $t$.
Remark 3.
In theorem 3, the naturally reductive structure does not come directly from $\mathrm{SO}(n)/\mathrm{SO}(n-d)$, but from $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$ with $\tilde{\mathcal{G}}=\mathrm{SO}(n)\times\mathrm{SO}(d)$ and $\tilde{\mathcal{K}}=\{(\operatorname{diag}(H,K),H)|H\in\mathrm{SO}(d),K\in\mathrm{SO}(n-d)\}$. We denote an element $\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$ of $\mathfrak{m}$ by $\mathtt{M}=\llceil A,B\rrceil$ with $A\in\mathfrak{so}(d),B\in\mathbb{R}^{(n-d)\times d}$, then $\widetilde{\mathtt{M}}=(\llceil 2\alpha A,B\rrceil,(2\alpha-1)A)\in\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}$ is the lift of $\mathtt{M}$. Set $\mathtt{M}_{i}=\llceil A_{i},B_{i}\rrceil$ for $i=1,2$ then $\mathcal{P}_{\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}}[\widetilde{\mathtt{M}}_{1},\widetilde{\mathtt{M}}_{2}]$ is equal to
$\begin{gathered}\mathcal{P}_{\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}}(\llceil 4\alpha^{2}[A_{1},A_{2}]-B_{1}^{\mathsf{T}}B_{2}+B_{2}^{\mathsf{T}}B_{1},2\alpha(B_{1}A_{2}-B_{2}A_{1})\rrceil,(2\alpha-1)^{2}[A_{1},A_{2}])\\ =\mathsf{lift}\llceil(4\alpha-1)[A_{1},A_{2}]-B_{1}^{\mathsf{T}}B_{2}+B_{2}^{\mathsf{T}}B_{1},2\alpha(B_{1}A_{2}-B_{2}A_{1})\rrceil=\mathsf{lift}\;\mathsf{AS}(\mathtt{M}_{1},\mathtt{M}_{2})\end{gathered}$
where we simplify $4\alpha^{2}-(2\alpha-1)^{2}=4\alpha-1$, and $\mathsf{AS}(\mathtt{M}_{1},\mathtt{M}_{2})\in\mathfrak{m}$ is defined by the condition above. We can verify eq. 9 directly for the metric $\alpha\operatorname{Tr}A^{\mathsf{T}}A+\operatorname{Tr}B^{\mathsf{T}}B$ from
$\langle\mathtt{M}_{1},\mathsf{AS}(\mathtt{M}_{2},\mathtt{M}_{3})\rangle_{\vec{\beta}}=\alpha\operatorname{Tr}((1-4\alpha)A_{1}[A_{2},A_{3}]+2A_{1}B_{2}^{\mathsf{T}}B_{3}+2A_{2}B_{3}^{T}B_{1}+2A_{3}B_{1}^{\mathsf{T}}B_{2})$
using $\operatorname{Tr}A_{1}B_{2}^{\mathsf{T}}B_{3}=\operatorname{Tr}(A_{1}B_{2}^{\mathsf{T}}B_{3})^{\mathsf{T}}=-\operatorname{Tr}A_{1}B_{3}^{\mathsf{T}}B_{2}$ and similar expressions. The right-hand side is cyclic invariant, implying eq. 9. See also [41, Theorem 2.1].
9. Flag manifolds with the canonical metric